# GMT: Guided Mask Transformer for Leaf Instance Segmentation

### v2

@@ 4.4 Ablation Study
Firstly, we assess different backbones with results shown in Table 3. We find that the smallest model (ResNet-50) achieves the best SBD and ||||DiC||||, while the largest model (SwinT-B) reaches the best BD. As discussed earlier at Section 4.2, better SBD and ||||DiC|||| but worse BD indicate that ResNet-50 more accurately identifies and segments all leaves, while SwinT-B can produce higher-quality masks for some instances. Additionally, due to the small size of plant datasets, larger backbones tend to overfit, which is why ResNet-50 is used for all other experiments.
Next, we evaluate each component of GMT (c.f. Section 3.2.2), to understand their contributions to the overall model performance. The results are shown in Table 4. Compared to the baseline (Mask2Former), we observe that using a single component or a combination of two tends to enhance specific aspects of the model’s abilities, but not all. For instance, using GPE alone (row (a)) slightly improves BD while maintaining SBD but worsens ||||DiC||||. However, when all three modules, GPE, GEFM, and GDPQ, are integrated, GMT outperforms Mask2Former across all metrics. This comprehensive improvement highlights the effectiveness of our overall design of combining all modules to address the complexities of leaf instance segmentation.
@@ 5 Conclusion
Given the complexity of plant structures and the typically small size of labelled datasets, leaf instance segmentation in plant images is challenging. To address this task, we present GMT, a Transformer-based model that effectively incorporates leaf position priors to better capture plant complexity on limited data. We demonstrate the effectiveness of GMT in three widely used datasets, showing that GMT outperforms the SotA, particularly in challenging cases such as smaller and overlapping leaves. Ablation studies further validate the backbone choice and overall design of GMT. Future work will focus on extending GMT to work on more complex scenarios such as in-field plant phenotyping and broader agricultural applications.

### v3

@@ 4.4 Ablation Study
Firstly, we assess different backbones with results shown in Tab. 3. We find that the smallest model (ResNet-50) achieves the best SBD and ||||DiC||||, while the largest model (SwinT-B) reaches the best BD. As discussed earlier at Sec. 4.2, better SBD and ||||DiC|||| but worse BD indicate that ResNet-50 more accurately identifies and segments all leaves, while SwinT-B can produce higher-quality masks for some instances. Additionally, due to the small size of plant datasets, larger backbones tend to overfit, which is why ResNet-50 is used for all other experiments.
Next, we evaluate each component of GMT (c.f. Sec. 3.2.2), to understand their contributions to the overall model performance. The results are shown in Tab. 4. Compared to the baseline (Mask2Former), we observe that using a single component or a combination of two tends to enhance specific aspects of the model’s abilities, but not all. For instance, using GPE alone (row (a)) slightly improves BD while maintaining SBD but worsens ||||DiC||||. However, when all three modules, GPE, GEFM, and GDPQ, are integrated, GMT outperforms Mask2Former across all metrics. This comprehensive improvement highlights the effectiveness of our overall design of combining all modules to address the complexities of leaf instance segmentation.
@@ 5 Conclusion
Given the complexity of plant structures and the typically small size of labelled datasets, leaf instance segmentation in plant images is challenging. To address this task, we present GMT, a Transformer-based model that effectively incorporates leaf position priors to better capture plant complexity on limited data. We demonstrate the effectiveness of GMT in three widely used datasets, showing that GMT outperforms the SOTA (Tab. 1), particularly in challenging cases such as smaller and overlapping leaves (Tab. 3 and Fig. 5). Ablation studies further validate the backbone choice and overall design of GMT. Future work will focus on extending GMT to address more complex agricultural scenarios while reducing its computational demands.
Acknowledgements
This project was funded by the BBSRC grant BB/Y512333/1 “PhenomUK-RI: The UK Plant and Crop Phenotyping Infrastructure”, and Microsoft Accelerating Foundation Models Research (AFMR) grant: Agricultural Foundation Models via Domain-Specific Pre-Training.